# Identifying and Fitting Eclipse Maps of Exoplanets with Cross-Validation

### v1

@@ 4.2 An optimised map fitted to observations of WASP-43b
Hammond et al. (2024) suggested that a uniform longitudinal structure near the substellar point could be caused by dayside stationary Rossby waves. These waves appear in some of the GCM simulations in Hammond et al. (2024) and produce a less sharply peaked longitudinal structure (Lewis & Hammond, 2022). However, they also produce a two-peaked latitudinal structure with maxima north and south of the equator (Matsuno, 1966), but the data are not sufficiently precise to resolve or rule out a two-peaked structure in Figure 14. More observations might constrain the dayside structure strongly enough to detect such structures.
@@ 5 Conclusions
We conclude that this method is an improvement on fitting eclipse maps using spherical harmonics alone without entropy-based regularisation. We showed that the method correctly identified the presence of an eclipse mapping signal in accurate and precise simulated datasets and the absence of an eclipse mapping signal in inaccurate or imprecise simulated datasets. We then fitted a map of appropriate precision to the datasets containing eclipse mapping signals.
@@ Appendix A Additional demonstration of optimising mapping information content
Figure 17: First row: an eclipse map model $\mathcal{M}(\ell_{\mathrm{max}}=4)$ fitted to a light curve simulated from Equation 16 with precision 150 ppm as described in Appendix A. Fitting the the spherical harmonics alone results in a highly imprecise map. Second row: an eclipse map model $\mathcal{M}(\alpha=4642,\ell_{\mathrm{max}}=4)$ with the parameter $\alpha$ optimised according to the cross-validation score shown in Figure 16.

### v2

@@ 4.2 An optimised map fitted to observations of WASP-43b
Hammond et al. (2024) suggested that a uniform longitudinal structure near the substellar point could be caused by dayside stationary Rossby waves. These waves appear in some of the GCM simulations in Hammond et al. (2024) and produce a less sharply peaked longitudinal structure (Lewis & Hammond, 2022). However, they also produce a two-peaked latitudinal structure with maxima north and south of the equator (Matsuno, 1966), but the data are not sufficiently precise to resolve or rule out a two-peaked structure in Figure 14. More observations might constrain the dayside structure strongly enough to detect such structures. We note that some of the the “observable” GCM maps in Hammond et al. (2024) have their two-peaked structures and their chevron-shaped structures suppressed compared to the original GCM simulation, making them more similar to the optimised map we derive here.
@@ 5 Conclusions
We conclude that this method is an improvement on fitting eclipse maps using spherical harmonics alone without entropy-based regularisation. Our approach is, we hope, a complementary method to the widely used eigenmapping fitting method (Rauscher et al., 2018). The eigenmapping method adjusts the number of functions used to fit the lightcurve to optimise a statistical metric, while our new method adjust the information content of the map to optimise a different statistical metric. We showed that our new method correctly identified the presence of an eclipse mapping signal in accurate and precise simulated datasets and the absence of an eclipse mapping signal in inaccurate or imprecise simulated datasets. We then fitted a map of appropriate precision to the datasets containing eclipse mapping signals.
@@ Appendix A Additional demonstration of optimising mapping information content
Figure 17: First row: an eclipse map model $\mathcal{M}(\ell_{\mathrm{max}}=4)$ fitted to a light curve simulated from Equation 16 with precision 150 ppm as described in Appendix A. The crosses on the maps show the position of the peak of the Gaussian perturbation defined in Equation 16. Fitting the the spherical harmonics alone results in a highly imprecise map. Second row: an eclipse map model $\mathcal{M}(\alpha=4642,\ell_{\mathrm{max}}=4)$ with the parameter $\alpha$ optimised according to the cross-validation score shown in Figure 16.